Circular No. 7904

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2002 J5 (LINEAR) An apparently asteroidal object discovered by LINEAR (discovery ob-

servation below) has been found to be slightly diffuse with a coma diameter of ∼ 10′′ on CCD images taken by M. Tichý and M. Kocer with the 1.06-m reflector at Kleť on May 16.0 UT.

2002 UT  α_{2000}  δ_{2000}  m_{2}
May 15.38149  20^{h}16^{m}07.^{s}57  +8^{o}44′51′′.5  18.2

The available astrometry (including prediscovery observations from Aug. 2001) and the following orbital elements appear on MPEC 2002-K16.

Epoch = 2003 Oct. 8.0 TT

T = 2003 Sept.19.31543 TT  ω = 74.^{o}83511
e = 1.0012361  Ω = 314.11014 }2000.0
q = 5.7266705 AU  i = 117.22838

COMET C/2002 K2 (LINEAR) Another apparently asteroidal LINEAR discovery has also been found

to be cometary on Kleť CCD images that show a compact coma of diameter ∼ 7′′–8′′ on May 16.9 and 18.0 UT.

2002 UT  α_{2000}  δ_{2000}  m_{2}
May 16.36948  18^{h}06^{m}11.^{s}35  +12^{o}03′22′′.6  18.2

The available astrometry and the following very uncertain parabolic orbital elements appear on MPEC 2002-K18.

T = 2002 Sept.29.407 TT  ω = 58.^{o}671
Ω = 296.557 }2000.0
q = 5.22393 AU  i = 132.541

COMET C/2002 K1 (NEAT) Additional astrometry (cf. IAUC 7902) and the following preliminary

parabolic orbital elements appear on MPEC 2002-K17.

T = 2002 Apr. 30.011 TT  ω = 353.^{o}116
Ω = 280.643 }2000.0
q = 3.24338 AU  i = 88.222
2002 May 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green